Use slopes to show that the square with vertices at (-2,5), (4,5), (4,-1), and (-2,-1)has diagonals that are perpendicular.
Answer by checkley71(8403)

SLOPE BETWEEN (-2,5) & (4,-1) IS (-1-5)/(4+2)=-6/6=-1
SLOPE BETWEEN (4,5) & (-2,-1) IS (-1-5)/(-2-4)=-6/-6=1
THESE TWO SLOPES ARE NEGATIVE RECIPRICALS OF EACH OTHER THUS THESE LINES ARE PERPENDICULAR.